OpenStudy (anonymous):

Need help with factoring this. x^2 - 8x + 15

OpenStudy (cwrw238):

these factor to the form ( x +/- a)( x =/- b) for your function you need 2 numbers a and b where a*b = + 15 and a + b = -8

OpenStudy (cwrw238):

* thats ( x +/- a)( x +/- b)

OpenStudy (anonymous):

x=3,5

OpenStudy (cwrw238):

factoring 15 will help = 3 * 5 or -3 * -5

OpenStudy (cwrw238):

the question is to factorize

OpenStudy (cwrw238):

-3 * -5 = 15 -3 + (-5) = ?

OpenStudy (anonymous):

Ok, I get it now . (x-3)(x-5)

OpenStudy (cwrw238):

right